A random sample of eight drugstores shows the following prices (in $ of a popular pain reliever:

3.50, 4.00, 2.00, 3.00, 2.50, 3.50, 2.50, 3.00

Assume the normal distribution for the underlying population to construct the 90% confidence interval for the population mean.



1
Expert's answer
2020-07-21T19:05:29-0400

"\text{The confidence interval for the mean }\\
\text{can be calculated as follows:}\\
 \bar x±t_{(\frac{α}{2},n-1)}(\frac{s}{\sqrt{n}}),\\
 \bar x=\frac{3.5+4+...+3}{8}=3\\
 t_{(\frac{α}{2},n-1)}=t_{(0.05,7)}=1.895,\\
 s^2=\frac{(3.5-3)^2+(4-3)^2+...+(3-3)^2}{7}=\frac{3}{7}\\
s=\sqrt{s^2}\approx0.655
\\n=8\\
 \text{ the lower limit }= 3-1.895(\frac{0.655}{\sqrt{8}})\\
≈2.56,\\
 \text{ the upper limit} = 3+1.895(\frac{0.655}{\sqrt{8}})\\
≈3.44,\\
 \text{so the confidence interval is}\\ 2.56\leq \bar x \leq 3.44\\"